We are to write the followingstatment using mathematical symbols:
**n is the product of the difference between n squared and seven, and n cubed**
I got the following two, I dont know which is correct or if either is. Can you please help me.
n^2-7(n^3)=n or n^2+7-n^3=n
Thank you so much!!!!

Answer by Edwin McCravy(20086)

n is the product of the difference between n squared and seven, and n cubed**
I got the following two, I dont know which is correct or if either is. Can you please help me.
n^2-7(n^3)=n or n^2+7-n^3=n
Thank you so much!!!!

Actually neither one is correct, though your first one is close.
Let's break it down.

n is the product of the difference between n squared and seven, and n cubed

Let's write " is " as "equals" or " = "

n = the product of the difference between n squared and seven, and n cubed

Let's write "n cubed" as " n³ "

n = the product of the difference between n squared and seven, and n³

Let's write "seven" as " 7 "

n = the product of the difference between n squared and 7, and n³

Let's write "n squared" as " n² ".

n = the product of the difference between n² and 7, and n³

Let's write "the difference between n² and 7" as " n² - 7 "

n = the product of n² - 7 and n³

Let's write "the product of n² - 7 and n³" as " (n² - 7)n³"

n = (n² - 7)n³

So your first answer would have been correct, if you had only
enclosed the n² - 7 in parentheses to get

(n² - 7)n³ = n

Edwin